### v1

# The special unitary groups $SU(2n)$ as framed manifolds

Abstract.
Let $[SU(2n),\mathscr{L}]$ denote the bordism class of $SU(2n)$ $(n\geq 2)$ equipped with the left invariant framing $\mathscr{L}$. Then it is well known that $e_{\mathbb{C}}([SU(2n),\mathscr{L}])=0$ in $\mathbb{O}/\mathbb{Z}$ where $e_{\mathbb{C}}$ denotes the complex Adams $e$-invariant. In this note we show that replacing $\mathscr{L}$ by the twisted framing by a specific map it can be transformed into a generator of $\mathrm{Im}\,e_{\mathbb{C}}$. In addition to that we also show that the same procedure affords an analogous result for a quotient of $SU(2n+1)$ by a circle subgroup which inherits a canonical framing from $SU(2n+1)$ in the usual way.
2020 Mathematics Subject Classification:
22E46, 55Q45
1. Introduction
Let $G$ be a simply-connected compact Lie group of dimension $14\mathit{l}-$ 1 and of rank $\geq 2$. Let us consider it to be the framed manifold equipped with the left invariant framing $\mathscr{L}$ and write $[G,\mathscr{L}]$ for its bordism class in $\pi_{4\mathit{l}-1}^{S}$. Then it is well known  [3] that
$e_{\mathbb{C}}([G,\mathscr{L}])=0$
where $e_{\mathbb{C}}:\pi_{4\mathit{l}-1}^{S}\to\mathbb{Q}/\mathbb{Z}$ denotes the complex Adams $e$-invariant. In view of this result we wish to raise the problem of the existence of a map $f:G\to GL(s,\mathbb{R})$ such that $e_{\mathbb{C}}([G,\mathscr{L}^{f}])$ is nonzero and if possible, is equal to $(-1)^{\ell-1}B_{\ell}/2\ell$  [2] where $\mathscr{L}^{f}$ denotes the framing obtained by twisting $\mathscr{L}$ by $f$ and $B_{\mathit{l}}$ the $\mathit{l}$th Bernoulli number.
In this note we consider the case $G=SU(2n)$ $(n\geq 2)$ where $\dim G=4n^{2}-1$. The main result is the following theorem.
Theorem .
Let $\rho:SU(2n)\to GL(4n,\mathbb{R})$ $(n\geq 2)$ be the standard real representation of $SU(2n)$. Then we have
$e_{\mathbb{C}}\bigl{(}[SU(2n),\mathscr{L}^{(n-1)\rho}]\bigr{)}=(-1)^{n-1}B_{n^{2}}/2n^{2}$
The proof is based on Proposition 2.1 of  [5]. But also in the case $G=SU(2n+1)$ it can be seen that the procedure used there is adaptable to its quotient by a circle subgroup $C$, considered as a framed manifold equipped with the induced framing $\mathscr{L}_{C}$. Now this framing can be obtained by applying the fact that $\mathscr{L}$ is a $C$-equivariant framing  [4] to the equation ()($\ast$)( ∗ ) of  [5, p. 16]. Here let $S^{1}=\{z\in\mathbb{C}||z|=1\}$ and put
$C=\{\mathrm{diag}(1,\ldots,1,z,\bar{z})\mid z\in S^{1}\}$
where $\mathrm{diag}(a_{1},\ldots,a_{2n+1})$ denotes the diagonal matrix with $a_{1},\ldots,a_{2n+1}$ on the diagonal. Then in fact we have
Proposition .
$e_{\mathbb{C}}\bigl{(}[SU(2n+1)/C,\mathscr{L}_{C}]\bigr{)}=-B_{n^{2}+n}/2(n^{2}+n)$  $(n\geq 1)$.
2. Tensor product decomposition of $E$
Let $G=SU(m)$ $(m\geq 2)$ and $S$ be the circle subgroup of $G$ generated by $d(z)=\mathrm{diag}(\bar{z}^{m-1},z,\ldots,z)$ with $z\in S^{1}$. Regard $G$ as an $S$-space endowed with $d(z)$ acting on $g\in G$ by the rule $(g,z)\to gd(z)$ and consider it to be the principal bundle along with the natural projection $p:G\to G/S$. Let $E=G\times_{S}\mathbb{C}$ denote the canonical complex line bundle over $G/S$ associated to $p$ where $S$ acts on $\mathbb{C}$ as $S^{1}$. Then its unit sphere bundle $\pi:S(E)\to G/S$ is naturally isomorphic to $p:G\to G/S$ as a principal $S$-bundle over $G/S$, which is expressed as $(S(E),\pi,G/S)\cong(G,p,G/S)$ as usual.
In order to introduce some further notations we first recall the case $m=2$. Write
$R(rz,v)=\begin{pmatrix}rz&v\\ -\bar{v}&r\bar{z}\end{pmatrix},\quad\ r\geq 0,\ z\in S^{1},\ v\in\mathbb{C},$
for the elements of $SU(2)$. Then $R(rz,v)d(z)=R(r,zv)$ and so letting it correspond to the element $(1-2r^{2},2rzv)\in S^{2}$ we have a homeomorphism between $SU(2)/S$ and $S^{2}$. Thinking of $p$ as a principal $S$-bundle over $S^{2}$ via this homeomorphism we have $(S(L),\pi,S^{2})\cong(SU(2),p,S^{2})$ where $L$ is used instead of $E$. From the formula of  [5, p. 40] we have $e([S(L),\Phi_{L}])=B_{1}/4$ where $\Phi_{L}$ denotes the trivialization of the stable tangent bundle of $S(L)$. In this case $\Phi_{L}$ coincides with $\mathscr{L}$ via the above isomorphism and so we have $e([SU(2),\mathscr{L}])=1/24$.
For brevity we write $(r,zv)_{R}$ for $(1-2r^{2},2rzv)\in S^{2}$ and identify
$(r,zv)_{R}=p(R(rz,v))=R(r,zv).$
Here in particular $(r,zv)_{R}$ represents $(0,1)_{R}$ if $r=0$ and $(1,0)_{R}$ if $r=1$. The latter is assumed to be specified as the base point of $S^{2}$ for the reason that it is the image of the identity matrix of $SU(2)$ by $p$.
Suppose $m\geq 3$ and let $20\leq j\leq m-$ 2 and $1\leq i\leq m_{j}$ where $m_{j}=m-j-1$. Further we assume that the factors used in the product constructions are arranged in ascending order of their suffixes derived by $i,j$. For fixed $j$ we put
$R_{i;j}(r_{i;j}z,v_{i;j})=\begin{pmatrix}I_{j}&0&0&0&0\\ 0&r_{i;j}z&0&v_{i;j}&0\\ 0&0&I_{i-1}&0&0\\ 0&-\bar{v}_{i;j}&0&r_{i;j}\bar{z}&0\\ 0&0&0&0&I_{m_{j}-i}\end{pmatrix}\ \text{with}\ R(r_{i;j}z,v_{i;j})\in SU(2)$
where $I_{s}$ denotes the identity matrix of order $s$. Then we have
(1) $\prod_{i=1}^{m_{j}}R_{i;j}(r_{i;j}z,v_{i;j})=\left(\begin{array}[]{cccccc}I_{j}&0&0&0&\cdots&0\\ 0&a_{0;j}&b_{1;j}&b_{2;j}&\cdots&b_{m_{j};j}\\ 0&a_{1;j}&c_{1,1;j}&c_{1,2;j}&\cdots&c_{1,m_{j};j}\\ 0&a_{2;j}&0&c_{2,2;j}&\cdots&c_{2,m_{j};j}\\ \vdots&\vdots&\vdots&\ddots&\ddots&\vdots\\ 0&a_{m_{j};j}&0&\cdots&0&c_{m_{j},m_{j};j}\end{array}\right)$
where
$\begin{split}&a_{0;j}=r_{1;j}\cdots r_{m_{j};j}z^{m_{j}},\\ &a_{s;j}=-r_{s+1;j}\cdots r_{m_{j};j}z^{m_{j}-s}\bar{v}_{s;j}\quad(1\leq s\leq m_{j}-1),\hskip 99.58464pt\\ &a_{m_{j};j}=-\bar{v}_{m_{j};j},\\ \end{split}$
[…]
Let $d_{j}(z)=\mathrm{diag}(1,\overset{(j)}{\ldots},1,\bar{z}^{m_{j}},z,\ldots,z)$. By referring to the above calculation we have
(2) $\biggl{(}\textstyle\prod_{i=1}^{m_{j}}R_{i;j}(r_{i;j}z,v_{i;j})\biggr{)}d_{j}(z)=\prod_{i=1}^{m_{j}}R_{i;j}(r_{i;j},z^{i}v_{i;j}).$
The matrix representation of the right-hand side of this equation has a similar form to that of (1). In fact, restricting ourselves to the part we need for later use we have
(3) $\begin{split}&\underline{a}_{0;j}=r_{1;j}\cdots r_{m_{j};j},\ \\ &\underline{a}_{s;j}=-r_{s+1;j}\cdots r_{m_{j};j}\bar{w}_{s;j}\ \ (1\leq s\leq m_{j}-1),\hskip 128.0374pt\\ &\underline{a}_{m_{j};j}=-\bar{w}_{m_{j};j},\\ \end{split}$
where $\underline{a}_{s;j}$ denotes its matrix element corresponding to $a_{s;j}$ and $w_{s;j}$ stands for $z^{s}v_{s;j}$.
Let $D_{0}(z)=I_{m}$ and $D_{j}(z)=\mathrm{diag}(z^{m-1},\bar{z},\overset{(j-1)}{\ldots},\bar{z},\bar{z}^{m-j},1,\ldots,1)$ $(1\leq j\leq m-2)$. Clearly $D_{j}(z)d(z)(=d(z)D_{j}(z))=d_{j}(z)$. Define $R^{\{i\}}_{j}(r_{i;j}z,v_{i;j})$ to be the product
$\biggl{(}\textstyle\prod_{s=1}^{m_{j}}R_{s;j}(r_{s;j}z,v_{s;j})\biggr{)}D_{j}(z)\ \text{with}\ r_{s;j}=1\ \text{for all}\ s,j\ \text{except}\ s=i.$
Then by (2) we have
(4) $R^{\{i\}}_{j}(r_{i;j}z,v_{i;j})d(z)=R_{i;j}(r_{i;j},z^{i}v_{i;j}),\qquad z\in S^{1}$
since $R_{s;j}(r_{s;j},z^{s}v_{sj})=I_{m}$ if $r_{s;j}=1$. Hence putting
$R_{j}(r_{i;j}z_{i},v_{i;j})=\textstyle\prod_{i=1}^{m_{j}}R^{\{i\}}_{j}(r_{i;j}z_{i},v_{i;j}),\qquad z_{i}\in S^{1}.$
we have due to the commutativeness of $d(z^{\prime})$ and $R^{\{i\}}_{j}(r_{i;j}z,v_{i;j})$ $(z,z^{\prime}\in S^{1})$
(5) $R_{j}(r_{i;j}z_{i},v_{i;j})d(z_{1})\cdots d(z_{m_{j}})=\textstyle\prod_{i=1}^{m_{j}}R_{i;j}(r_{i;j},z_{i}^{i}v_{i;j})\qquad(j\geq 1).$
But in the case of $j=0$, differing from this, we have
(5b) $R_{0}(r_{i;0}z_{i},v_{i;0})d(z_{1})\cdots d(z_{m-1})=R_{1;0}(r_{1;0},z_{1}v_{1;0})\textstyle\prod_{i=2}^{m-1}R_{i;0}(r_{i;0},z_{i-1}^{m}z_{i}^{i}v_{i;0}).$
This can be shown using induction. In fact, as in (5), we first have
$R_{0}(r_{i;0}z_{i},v_{i;0})d(z_{m-1})=\biggl{(}\textstyle\prod_{s=1}^{m-2}R_{i;0}(r_{i;0}z_{i},v_{i;0})\biggr{)}R_{m-1;0}(r_{m-1;0},z_{m-1}^{m-1}v_{m-1;0}).$
By letting $d(z_{m-2})$ act on both sides of this it can be derived that
$\begin{split}R_{0}(r_{i;0}z_{i},&v_{i;0})d(z_{m-1})d(z_{m-2})=\biggl{(}\textstyle\prod_{i=1}^{m-3}R_{i;0}(r_{i;0}z_{i},v_{i;0})\biggr{)}\\ &R_{m-2;0}(r_{m-2;0},z_{m-2}^{m-2}v_{m-2;0})R_{m-1;0}(r_{m-1;0},z_{m-2}^{m}z_{m-1}^{m-1}v_{m-1;0}).\end{split}$
Repeating this procedure we can achieve the equation (5b). Further these equations have the following extensions.
(6) $\begin{split}R^{\{i\}}_{j}(r_{i;j}zz^{\prime},v_{i;j})d(z)&=R^{\{i\}}_{j}(r_{i;j}z^{\prime}_{,}z^{i}v_{i;j})\quad(j\geq 0),\\ R_{j}(r_{i;j}z_{i}z^{\prime}_{i},v_{i;j})d(z_{1})\cdots d(z_{m_{j}})&=\begin{cases}R_{j}(r_{i;j}z^{\prime}_{s},z_{i}^{i}v_{i;j})&(j\geq 1),\\ R_{j}(r_{i;j}z^{\prime}_{i},z_{i-1}^{n}z_{i}^{i}v_{i;j})&(j=0)\quad(\text{where}\ z_{-1}=1).\end{cases}\end{split}$
From (4) we see that $p(R^{\{i\}}_{j}(r_{i;j}z,v_{i;j}))=p(R_{i;j}(r_{i;j},z^{i}v_{i;j}))$ holds for $p:G\to G/S$. This permits us to embed $S^{2}_{i;j}$ in $G/S$ by means of the injective map $(r_{i;j},w_{i;j})_{R}\to p(R^{\{i\}}_{j}(r_{i;j}z,v_{i;j}))$ where $S^{2}_{i;j}=S^{2}$ denotes the 2-sphere consisting of $(r_{i;j},w_{i;j})_{R}$. From (6) we also see that
$P^{\{i\}}_{j}=\{R^{\{i\}}_{j}(r_{i;j}z,v_{i;j})|(r_{i;j},w_{i;j})_{R}\in S^{2},z\in S^{1}|\}\subset G$
is closed under the action of $S$. Combining these two facts we have that $q_{i;j}=p|P^{\{i\}}_{j}:P^{\{i\}}_{j}\to S^{2}_{i;j}\subset G/S$ provides the projection map of a principal $S$-bundle over $S^{2}_{i;j}$. Then $q_{i;j}$ factors through the map $P^{\{i\}}_{j}\to SU(2)$ given by $R^{\{i\}}_{j}(r_{i;j}z,v_{i;j})\to R(r_{i;j}z,z^{i-1}v_{i;j})$ which yields an isomorphism $(P^{\{i\}}_{j},q_{i;j},S^{2}_{i;j})\cong(SU(2),p,S^{2})$. Let us write $L_{i;j}$ for the canonical complex line bundle over $S^{2}_{i;j}$ associated to $q_{i;j}$. Then since $(S(L_{i;j}),\pi,S^{2}_{i;j})\cong(P^{\{i\}}_{j},q_{i;j},S^{2}_{i;j})$ we have
(7) $(S(L_{i;j}),\pi,S^{2}_{i;j})\cong(SU(2),p,S^{2})\quad\ (\text{identifying}\ S^{2}_{i;j}=S^{2}).$
Let us write $(S_{j}^{2})^{m_{j}}=S^{2}_{1;j}\times\cdots\times S^{2}_{m_{j};j}$ and let $\phi_{j}:(S_{j}^{2})^{m_{j}}\to G/S$ be the map given by $x_{j}=(x_{1;j},\cdots,x_{m_{j};j})\to p(R_{j}(r_{i;j},w_{i;j}))$ where $x_{i;j}=(r_{i;j},w_{i;j})_{R}\in S^{2}_{i;j}$. Let us also write $(S^{2})^{(m^{2}-m)/2}=(S_{0}^{2})^{m_{0}}\times\cdots\times(S_{m-2}^{2})^{m_{m-2}}$ and let $\phi:(S^{2})^{(m^{2}-m)/2}\to G/S$ be the map defined by $\phi(x)=\phi_{0}(x_{0})\cdots\phi_{m-2}(x_{m-2})$ where $x=(x_{0},\cdots,x_{m-2})$ with $x_{j}\in(S_{j}^{2})^{m_{j}}$. Let us put $R(r_{i;j}z_{i},v_{i;j})=\prod_{j=0}^{m-2}R_{j}(r_{i;j}z_{i},v_{i;j})$. Then, as in the case of $P^{\{i\}}_{j}$, using (5), (5b), (6),
$P=\{R(r_{i;j}z_{i},v_{i;j})|(r_{i;j},w_{i;j})_{R}\in S^{2}_{i;j},z_{i}\in S^{1}\}\subset G$
forms the total space of a principal $S$-bundle endowed with the projection map $q:P\to(S^{2})^{(m^{2}-m)/2}$ such that $\phi\circ q=p|P$. Let us write $L_{j}^{m_{j}}=L_{1;j}\boxtimes\cdots\boxtimes L_{m_{j};j}$ and also put $L^{(m^{2}-m)/2}=L_{0}^{m_{0}}\boxtimes\cdots\boxtimes L_{m-2}^{m_{m-2}}$. Then $(S(L^{(m^{2}-m)/2}),\pi,(S^{2})^{(m^{2}-m)/2})\cong(P,q,(S^{2})^{(m^{2}-m)/2})$ for the reason similar to (7) and therfore we have
(8) $(S(L^{(m^{2}-m)/2}),\pi,(S^{2})^{(m^{2}-m)/2})\cong\phi^{*}(G,p,G/S)$
where the right hand side denotes the induced bundle of $(G,p,G/S)$ by $\phi$.
Lemma 1.
$\phi:(S^{2})^{(m^{2}-m)/2}\to G/S$ is injective.
Proof.
By assumption $\phi(x)=\phi_{0}(x_{0})\cdots\phi_{m-2}(x_{m-2})$ and $\phi_{j}(x)=p(R_{j}(r_{i;j},w_{i;j}))$. Here $R_{j}(r_{i;j},w_{i;j})$ can be written in the form $R_{j}(r_{i;j},w_{i;j})=\mathrm{diag}(I_{j},M_{j})$ where the first column of $M_{j}$ consists of $\underline{a}_{0;j}$, $\underline{a}_{1;j}$, $\ldots$, $\underline{a}_{m-j-1;j}$ given in (3).
Suppose $\phi(x)=\phi(x^{\prime})$, that is, $\phi_{0}(x_{0})\cdots\phi_{m-2}(x_{m-2})=\phi_{0}(x^{\prime}_{0})\cdots\phi_{m-2}(x^{\prime}_{m-2})$. Then due to the definition this can be interpreted as meaning that
$R_{0}(r_{i;0},w_{i;0j})\cdots R_{m-2}(r_{i;m-2},w_{i;m-2})=R_{0}(r^{\prime}_{i;0},w^{\prime}_{i;0j})\cdots R_{m-2}(r^{\prime}_{i;m-2},w^{\prime}_{i;m-2})$
where we denote by attaching to an element accompanied by $x$ its corresponding element accompanied by $x^{\prime}$. In view of the form of $R_{j}(r_{i;j},w_{i;j})$ we therefore have
$\underline{a}_{0;0}=\underline{a}^{\prime}_{0;0},\ldots,\underline{a}_{m-1;0}=\underline{a}^{\prime}_{m-1;0}.$
Substituting the result of (3) into these equations we have $r_{1;0}\cdots r_{m-1;0}=r^{\prime}_{1;0}\cdots r^{\prime}_{m-1;0}$, $r_{s+1;0}\cdots r_{m-1;0}\bar{w}_{s;0}=r^{\prime}_{s+1;0}\cdots r^{\prime}_{m-1;0}\bar{w}^{\prime}_{s;0}$ $(1\leq s\leq m-1)$ and $\bar{w}_{m-1;0}=\bar{w}^{\prime}_{m-1;0}$ where $r_{s;0}^{2}+|w_{s;0}|^{2}=1$ and $r_{s;0}\geq 0$. For convenience we assume that $r_{i;0}>0$ and moreover proceed by induction on $i$ in reverse order. First from $\underline{a}_{m-1;0}=\underline{a}^{\prime}_{m-1;0}$ we have $w_{m-1;0}=w^{\prime}_{m-1;0}$ and so it follows that $r_{m-1;0}=r^{\prime}_{m-1;0}$. Next from the second equation $\underline{a}_{m-2;0}=\underline{a}^{\prime}_{m-2;0}$ we see that $r_{m-1;0}w_{m-2;0}=r^{\prime}_{m-1;0}w^{\prime}_{m-2;0}$. But since $r_{m-1;0}=r^{\prime}_{m-1;0}>0$ we have $w_{m-2;0}=w^{\prime}_{m-2;0}$ and so similarly it follows that $r_{m-2;0}=r^{\prime}_{m-2;0}$. Repeating this procedure we obtain subsequently $w_{m-3;0}=w^{\prime}_{m-3;0}$, $r_{m-3;0}=r^{\prime}_{m-3;0}$, $w_{m-4;0}=w^{\prime}_{m-4;0}$, $r_{m-4;0}=r^{\prime}_{m-4;0}$$\cdots$, $r_{2;0}=r^{\prime}_{2;0}$ and finally we obtain $w_{1;0}=w^{\prime}_{1;0}$, $r_{1;0}=r^{\prime}_{1;0}$. At this point we conclude that $x_{0}=x^{\prime}_{0}$ and so the assumption above can be rewritten as
$\phi_{1}(x_{1})\cdots\phi_{m-2}(x_{m-2})=\phi_{1}(x^{\prime}_{1})\cdots\phi_{m-2}(x^{\prime}_{m-2}).$
From this by looking at the form of $R_{j}(r_{i;j},w_{i;j})$ again we have $\underline{a}_{0;1}=\underline{a}^{\prime}_{0;1},\ldots,\underline{a}_{m-2;1}=\underline{a}^{\prime}_{m-2;1}$. Replacing the previous equations by these ones and the condition $r_{i;0}>0$ by $r_{i;1}>0$ in the argument there we can get $x_{1}=x^{\prime}_{1}$, hence the rewritten assumption above is also rewritten as
$\phi_{2}(x_{2})\cdots\phi_{m-2}(x_{m-2})=\phi_{2}(x^{\prime}_{2})\cdots\phi_{m-2}(x^{\prime}_{m-2}).$
Repeating this step we obtain $x_{2}=x^{\prime}_{2}$, $\ldots$, $x_{m-2}=x^{\prime}_{m-2}$ via inductive reduction, which implies $x=x^{\prime}$. This proves the lemma.
In the above, if $r_{m_{j};j}=0$, then we think of it as not being zero by exchanging the $m_{j}$th and other proper rows and columns depending on the case. In the case that there are no such row and column, this approach leads us to obtain directly $r_{i;j}=r^{\prime}_{i;j}=0$ $(1\leq i\leq m_{j})$ and so $x_{j}=x^{\prime}_{j}$. ∎
3. Proofs of Theorem and Proposition
For any $k\in\mathbb{Z}$ such that $11\leq k\leq m/2-$ 1 we put
$D(z_{2k-1},zz_{2k})=\mathrm{diag}(1,\overset{(2k-1)}{\ldots},1,z_{2k-1},\bar{z}_{2k-1}zz_{2k},\bar{z}\bar{z}_{2k},1,\ldots,1)$
for $z_{2k-1},z_{2k},z\in S^{1}$. Let $Q_{k}$ be the subspace of $G$ consisting of $D(z_{2k-1},zz_{2k})d(\bar{z})$. Then it forms the total space of a principal $S$-bundle over $T_{k}=S^{1}\times S^{1}$ along with the projective map $p_{k}:Q_{k}\to T_{k}$ given by $D(z_{2k-1},zz_{2k})d(\bar{z})\to(z_{2k-1},zz_{2k})$ where $T_{k}$ is considered as an subspace of $G/S$ under $\psi_{k}:(z_{2k-1},zz_{2k})\to p(D(z_{2k-1},zz_{2k}))$.
Let $\mu:T_{k}\to S^{2}$ be the map defined by $\mu(e^{\eta i},ze^{\theta i})=(-\cos\mu,ze^{\theta i}\sin\mu)$ where $z_{2k-1}=e^{\eta i}$ $(0\leq\eta\leq\pi)$ and $z_{2k}=e^{\theta i}$ $(0\leq\theta<2\pi)$. Taking into account the fact that a principal circle bundle over $S^{1}$ is trivial we see that the classifying map of $p_{k}$ factors through $S^{2}$ where the restriction of $p_{k}$ to $\{1\}\times S^{1}\subset T_{k}$ is viewed as being trivial. Then we have
Lemma 2.
$(Q_{k},p_{k},T_{k})\cong\mu^{*}(SU(2),p,S^{2})$  $(1\leq k\leq m/2-1)$.
Proof.
In order to prove this it suffices to show that there is a bundle map covering $\mu$. From the fact that $\mu$ can be rewitten as $\mu(e^{\eta i},ze^{\theta i})=(\cos(\eta/2),ze^{\theta i}\sin(\eta/2))_{R}$ we see that the map $\tilde{\mu}:Q_{k}\to SU(2)$ given by $D(e^{\eta i},ze^{\theta i})d(\bar{z})\to(\bar{z}\cos(\eta/2),e^{\theta i}\sin(\eta/2))$ in fact provides the required bundle map. ∎
Let $m=2n$ or $12n+$ 1 and put $T^{n-1}=T_{1}\times\cdots\times T_{n-1}$. Let $\phi_{j}:x_{j}\to p(R_{j}(r_{i;j},w_{i;j}))$ and $\psi_{k}:y_{k}\to p(D(z_{2k-1},\zeta_{2k}))$ be the injective maps of $(S_{j}^{2})^{m_{j}}$ and $T_{k}$ into $G/S$, respectively, described above where $x_{j}=(x_{1;j},\cdots,x_{m_{j};j})$ with $x_{i;j}=(r_{i;j},w_{i;j})_{R}\in S^{2}_{i;j}$ and $y_{k}=(z_{2k-1},\zeta_{2k})\in T_{k}$. Then by putting for $x=(x_{0},\cdots,x_{m-2})$ and $y=(y_{1},\cdots,y_{n-1})$
$\psi(x,y)=\phi_{0}(x_{0})\cdots\phi_{m-2}(x_{m-2})\psi_{1}(y_{1})\cdots\psi_{n-1}(y_{n-1})$
we have a map of $(S^{2})^{(m^{2}-m)/2}\times T^{n-1}$ into $G/S$. Let us put
$\tilde{R}(r_{i;j}z_{i},v_{i;j};z_{2k-1},z^{\prime}_{k}z_{2k})=\textstyle\prod_{j=0}^{m-2}R_{j}(r_{i;j}z_{i},v_{i;j})\prod_{k=1}^{n-1}D(z_{2k-1},z^{\prime}_{k}z_{2k})d(\bar{z}^{\prime}_{k}).$
Then as in the case of $P$
$\tilde{P}=\{\tilde{R}(r_{i;j}z_{i},v_{i;j};z_{2k-1},z^{\prime}_{k}z_{2k})|(r_{i;j},w_{i;j})_{R}\in S^{2}_{i;j},z^{\prime}_{k},z_{2k-1},z_{2k}\in S^{1}\}\subset G$
forms the total space of a principal $S$-bundle endowed with the projection map $\tilde{q}:\tilde{P}\to(S^{2})^{(m^{2}-m)/2}\times T^{n-1}$ such that $\psi\circ\tilde{q}=p|P$. Let $\tilde{L}_{k}$ be the complex line bundle over $T_{k}$ associated to $p_{k}$. Write $\tilde{L}^{n-1}=\tilde{L}_{1}\boxtimes\cdots\boxtimes\tilde{L}_{n-1}$ and put $E=L^{(m^{2}-m)/2}\boxtimes\tilde{L}^{n-1}$. Then we have $(S(E),\pi,(S^{2})^{(m^{2}-m)/2}\times T^{n-1})\cong(\tilde{P},q,(S^{2})^{(m^{2}-m)/2}\times T^{n-1})$ and therefore
(9) $(S(E),\pi,(S^{2})^{(m^{2}-m)/2}\times T^{n-1})\cong\psi^{*}(G,p,G/S).$
Lemma 3.
$\psi:(S^{2})^{(m^{2}-m)/2}\times T^{n-1}\to G/S$ is injective.
Proof.
In terms of the notation of the proof of Lemma 1 we prove that if we suppose that $\psi(x,y)=\psi(x^{\prime},y^{\prime})$, then it follows that $(r_{i;j},w_{i;j})_{R}=(r^{\prime}_{i;j},w^{\prime}_{i;j})_{R}$ $(0\leq j\leq m-2)$ and $(z_{2k-1},\zeta_{2k})=(z^{\prime}_{2k-1},\zeta^{\prime}_{2k})$ $(1\leq k\leq n-1)$, the latter of which is the part added to the proof of injectivity of $\phi$. The proof can be proceeded along the same lines as in the proof of the injectivity of $\phi$ above. But in fact there are the following changes in the use of the elements $\bar{w}_{m_{j};j}$ each of which is a starting point of the proof for $\phi_{j}$: $\bar{w}_{m-1;0}\to\bar{w}_{m-1;0}$ (i.e. no change), $\bar{w}_{m-2:1}\to z_{1}\bar{w}_{m-2:1}$, $\bar{w}_{m-3:2}\to(\bar{z}_{1}\zeta_{2})\bar{w}_{m-3:2}$ and in general $\bar{w}_{m-2k-2;2k+1}\to(z_{2k+1}\bar{\zeta}_{2k})\bar{w}_{m-2k-2;2k+1}$, $\bar{w}_{m-2k-3;2k+2}\to(\bar{z}_{2k+1}\zeta_{2k+2})\bar{w}_{m-2k-3;2k+2}$ for $k=1,\ldots,n-2$. Similarly substituting these elements into the equations $\underline{a}_{0;j}=\underline{a^{\prime}}_{0;j}$, $\underline{a}_{1;j}=\underline{a^{\prime}}_{1;j}$, $\ldots$, $\underline{a}_{m-j-1;j}=\underline{a^{\prime}}_{m-j-1;j}$ we can obtain $(x,y)=(x^{\prime},y^{\prime})$.
First from $\bar{w}_{m-1;0}=\bar{w}^{\prime}_{m-1;0}$ we have $x_{0}=x^{\prime}_{0}$, which shows that $z_{1}\bar{w}_{m-2:1}=z^{\prime}_{1}\bar{w}^{\prime}_{m-2:1}$, so we have $z_{1}r_{m-2:1}=z^{\prime}_{1}r^{\prime}_{m-2:1}$. If $r_{m-2:1}>0$, then from the second equation it follows that $r_{m-2:1}=r^{\prime}_{m-2:1}$ and so $z_{1}=z^{\prime}_{1}$. If $r_{m-2:1}=0$, then since $x_{0}=x^{\prime}_{0}=(0,1)_{R}$ are satisfied we get $z_{1}=z^{\prime}_{1}$. However in this case in order to move the next step we need to use the approrch of exchanging rows and columns. In any case it can be concluded that $x_{1}=x^{\prime}_{1}$. $z_{1}=z^{\prime}_{1}$, Next from $(\bar{z}_{1}\zeta_{2})\bar{w}_{m-3:2}=(\bar{z}^{\prime}_{1}\zeta^{\prime}_{2})\bar{w}^{\prime}_{m-3:2}$ we have $\zeta_{2}\bar{w}_{m-3:2}=\zeta^{\prime}_{2}\bar{w}^{\prime}_{m-3:2}$ and so it follows that $x_{2}=x^{\prime}_{2}$, $\zeta_{2}=\zeta^{\prime}_{2}$. Therefore by $(z_{3}\bar{\zeta}_{2})\bar{w}_{m-4;3}=(z^{\prime}_{3}\bar{\zeta}^{\prime}_{2})\bar{w}^{\prime}_{m-4;3}$ we have $z_{3}\bar{w}_{m-4;3}=z^{\prime}_{3}\bar{w}^{\prime}_{m-4;3}$ and so in a similar way to the previous case we have $x_{3}=x^{\prime}_{3}$, $z_{3}=z^{\prime}_{3}$. Subsequently also by $(\bar{z}_{3}\zeta_{4})\bar{w}_{m-5;4}=(\bar{z}^{\prime}_{3}\zeta^{\prime}_{4})\bar{w}^{\prime}_{m-5;4}$ we have $\zeta_{4}\bar{w}_{m-5;4}=\zeta^{\prime}_{4}\bar{w}^{\prime}_{m-5;4}$ and so it follows that $x_{4}=x^{\prime}_{4}$, $\zeta_{4}=\zeta^{\prime}_{4}$. At this point we can conclude that $x_{1}=x^{\prime}_{1}$, $x_{2}=x^{\prime}_{2}$, $x_{3}=x^{\prime}_{3}$, $x_{4}=x^{\prime}_{4}$ and $y_{1}=y^{\prime}_{1}$, $y_{2}=y^{\prime}_{2}$. These tell us that through repetition of these procedures we can arrive at the desired result. ∎
Proof of Theorem.
First we note that $\dim((S^{2})^{(m^{2}-m)/2}\times T^{n-1})=\dim G/S$. In general $SU(m)$ is embedded in $\mathbb{C}^{m(m-1)}$ as the subspace of $(m-1)$-tuples $(x_{1},\cdots,x_{m-1})$ with $x_{i}\in\mathbb{C}^{m}$, $|x_{i}|=1$ and $x_{i}\bot x_{j}$ if $i\neq j$ (cf.  [1] and  [6]). In view of this, if we think of $\psi$ as a map between manifolds of the same dimension $m^{2}-2$, then by taking into account its differentiability we can see from Lemma 3 that $\psi$ is a homeomorphism and so by construction it can also be taken to be a diffeomorphism.
Identify $G/S$ with $(S^{2})^{(m^{2}-m)/2}\times T^{n-1}$ via $\psi$ and write $E$ for $L^{(m^{2}-m)/2}\boxtimes\tilde{L}^{n-1}$ in (9). Then combining the equations of (7) and Lemma 2 we have from Proposition 2.1 of [5]
$e_{\mathbb{C}}([S(E),\Phi_{E}])=(-1)^{n-1}B_{n^{2}}/2n^{2}$
where $\Phi_{E}$ denotes the trivialization of the stable tangent space of $S(E)$ derived by the induced framing $\mathscr{L}_{S}$ on $G/S$.
Now $D(z_{2k-1},zz_{2k})\in Q_{k}$ equals $D(z_{2k-1},zz_{2k})=R_{1;2k-1}(z_{2k-1},0)R_{1;2k}(zz_{2k},0)$ where $R_{1;2k-1}(z_{2k-1},0)$ and $R_{1;2k}(zz_{2k},0)$, respectively, constitute the product components of $P\subset G$ which belong to $q_{0,2k-1}^{-1}((1,0)_{R})$ and $q_{0,2k}^{-1}((1,0)_{R})$. Putting $\mathscr{L}$ on $\Phi_{E}$, from this we see that the restrictions of $\mathscr{L}$ to $Q_{k}$ and $P$ overlap and therefore the restriction of $\mathscr{L}$ to $T_{k}\subset G$ must be stably trivial. This tells us that if we let $\rho$ denote the realification of the standard complex representation $G\hookrightarrow GL(m,\mathbb{C})$, then $\Phi_{E}$ must be stably equivalent to $\mathscr{L}^{(n-1)\rho}$. Substituting this into the equation above we have the desired result. ∎
Proof of Proposition.
Let $m=2n+1$. Let $C$ be as above. Then clearly $S\cap C=\{I_{2n+1}\}$, so $S\times C$ becomes a subgroup of $G$. Letting $r:G\to G/C$ and $\tilde{r}:G/C\to G/(S\times C)$ be the quotient maps, the pair $(r,\tilde{r})$ gives a bundle map between principal $S$-bundles $G\to G/S$ and $G/C\to G/(S\times C)$. Consider the composition $r\circ\iota:\tilde{P}\hookrightarrow G\to G/C$ and $\tilde{r}\circ\psi:(S^{2})^{(m^{2}-m)/2}\times T^{n-1}\to G/S\to G/(S\times C)$ where $\iota$ denotes the inclusion of $\tilde{P}$ into $G$. Then clearly $(\tilde{r}\circ\psi)\circ\tilde{q}=\tilde{r}\circ(r\circ\iota)$ holds and $\dim((S^{2})^{(m^{2}-m)/2}\times T^{n-1})=\dim G/(S\times C)$. Following the proof of Lemma 3 we can see that $\tilde{r}\circ\psi$ is injective and so it becomes a homeomorphism. Therefore by replacing $\psi$ by $\tilde{r}\circ\psi$ in the proof of the theorem above we can get a modified form in the case of $m=2n+1$. This proves the proposition. ∎